A spring vibrates with of frequency of 3 Hz when a mass of 1 kg is hung from it. Find the spring costant K. If a 3 kg mass were added to the 1kg mass, what would be the resulting frequency?
Answer #1

a)

w ^2 = k/m

( 2 *3.14* 3)^2 = k/ 1

k = 354.95 N/m

=======

b)

(2 pi f) ^2 = 354.95 / 4

f = 1.5 Hz

========

Comment in case any doubt, will reply for sure.. Goodluck